Show that the sequence (an), where an= n/(n^2+4) is monotonic. Is (an ) a Cauchy sequence? Justify your answer


1
Expert's answer
2021-04-13T13:57:00-0400

Let us show that the sequence "(a_n)", where "a_n= \frac{n}{n^2+4}" is monotonic. For this let us find the difference


"a_{n+1}-a_n=\frac{n+1}{(n+1)^2+4}-\frac{n}{n^2+4}=\frac{(n+1)(n^2+4)-n((n+1)^2+4)}{n^2+4}=
\frac{n^3+n^2+4n+4-n(n^2+2n+5)}{n^2+4}=\frac{-n^2-n+4}{n^2+4}<0" for "n\ge2."


It folllows that "a_{n+1}<a_n" for "n\ge 2", and hence the sequence is decreasing for "n\ge 2."


Let us show that "(a_n)" is a Cauchy sequence. For any "\varepsilon>0" let "n=\lceil\frac{2}{\varepsilon}\rceil". Then for any "m\ge n,\ k\ge n" we have that "|a_m-a_k|\le |a_m|+|a_k|=\frac{m}{m^2+4}+\frac{k}{k^2+4}<\frac{m}{m^2}+\frac{k}{k^2}=\frac{1}{m}+\frac{1}{k}\le\frac{2}{n}\le\varepsilon."

Therefore, for any "\varepsilon>0" there exists "n\in\mathbb N" such that for any "m\ge n,\ k\ge n" we have that "|a_m-a_k|<\varepsilon", and we conclude that "(a_n)" is a Cauchy sequence.